OpenStudy (anonymous):

solve cosx + sinx = cos2x +sin2x

OpenStudy (xapproachesinfinity):

So what do plan to do, what are you thoughts to start solving this?

OpenStudy (anonymous):

simplify cos2x and sin2x ?

OpenStudy (xapproachesinfinity):

For example cos2x is what, sin2x is what?

OpenStudy (xapproachesinfinity):

Okay try that!

OpenStudy (anonymous):

tried...still im not getting it...

OpenStudy (xapproachesinfinity):

show me what did you get!

OpenStudy (anonymous):

cosx + sinx = (cosx - sinx) * ( cosx + sinx) + 2sinxcosx

OpenStudy (anonymous):

correct up to this ?

OpenStudy (xapproachesinfinity):

Yes! what else did you do?

OpenStudy (xapproachesinfinity):

simplify those parenthesis, you could find new leads!

OpenStudy (lochana):

cosx + sinx = cos2x + sin2x sin2x - sinx = cos2x - cosx 2cos(3x/2)sin(x/2) = 2sin(3x/2)sin(-x/2) cos(3x/2)= -sin(3x/2) tan(3x/2)=-1

OpenStudy (lochana):

oh i forgot something, tan(3x/2) = tan(3pi/4) therefor , 3x/2 = n(pi) + 3(pi)/4 x =2(pi)(n+ 3/4)/3

OpenStudy (xapproachesinfinity):

from here cosx + sinx = cos2x + sin2x to here sin2x - sinx = cos2x - cosx a sign check the work again

OpenStudy (xapproachesinfinity):

And please explain your steps!

OpenStudy (anonymous):

thankyou so much lochana...

OpenStudy (anonymous):

my steps were a little complicated i guess...i used the formula of cos2x and sin2x directly cos2x = cos^2x - sin^2x and sin2x = 2cosx*sinx

OpenStudy (anonymous):

and thankyou xapproachesinfinity for giving me some sense

OpenStudy (xapproachesinfinity):

Anytime! there are many ways to solving trig equations! you just need to see what works with what!

OpenStudy (xapproachesinfinity):

@lochana please next time explain your steps!

OpenStudy (anonymous):

yeah, thats where the trouble comes in. i dont know when to use which formula...so it all just jumbles up into a mess leaving it more complicated than it is... but thanks anyways :)

OpenStudy (xapproachesinfinity):

It's okay, we all share that! we all struggle in the beginning, doing this continuously will help you recognize your mistakes and direct you to find or solving problems a lot easier it is a habit by training your mind to this thin

OpenStudy (xapproachesinfinity):

And what I got tan(3x/2)=1 not negative one

OpenStudy (xapproachesinfinity):

I was pushing you to do it and see where you struggle and how you would think it over haha but lochana intervened lol

OpenStudy (lochana):

@xapproachesinfinity yes you are correct, tan(3x/2)=1,

OpenStudy (xapproachesinfinity):

Alright! thanks!